OpenStudy (anonymous):

solve the system any algebraic method. 3.3x-1.5y=5.22 1.1x+2.6=7.32

hero (hero):

y = (5.22 - 3.3x)/1.5 y = (7.32-1.1x)/2.6 Set y = y (5.22 - 3.3x)/1.5 = (7.32-1.1x)/2.6 Cross Multiply: 13.572-8.58x = 10.98 - 1.65x 6.93x = 2.592 x = 0.374026

hero (hero):

y = 2.66

OpenStudy (dumbcow):

that works you could also multiply by 100 to get rid of decimals

hero (hero):

Yeah, I didn't think to do that

hero (hero):

For some reason, I just like using the original numbers

hero (hero):

Either way, you end up with a decimal answer